Algebraically determine the EXACT roots in simplest form for:
a) (5/x-3)=(x/x+1)
b) 3x-2=(-13/x+4)

Answer by stanbon(75887)

Algebraically determine the EXACT roots in simplest form for:
a) (5/x-3)=(x/x+1)
Cross-multiply to get:
5(x+1) = x(x-3)
5x+5 = x^2-3x
x^2-8x-5 = 0
x = [8 +- sqrt(64-4*-5)]/2
x = [8 +- sqrt(84)]/2
x = [8 +- 2sqrt(21)]/2
x = 4 +- sqrt(21)
========================
b) 3x-2=(-13/x+4)
(3x-2)(x+4)+13 = 0
3x^2+10x+5 = 0
x= [-10 +- sqrt(100-4*3*5)]/6
x= [-10 +- sqrt(40)]/6
x = [-10 +- 2sqrt(10)]/6
x = [(-5/3) +- (1/3)sqrt(10)]
==================================
Cheers,
Stan H.